how many liters of a 8% solution of salt should be added to a 23% solution to obtain 645 liters of a 14% solution
Answer by lwsshak3(11628)

how many liters of a 8% solution of salt should be added to a 23% solution to obtain 645 liters of a 14% solution
***
let x=amt of 8% solution to be added
645-x=amt of 23% solution to be used
..
8%x+23%(645-x)=14%*645
.08x+148.35-.23x=90.3
.15x=50.05
x=387
liters of a 8% solution of salt to be added=387